OpenStudy (anonymous):

Which equation has intercepts X(1, 0, 0), Y(0, –1, 0), and Z(0, 0, 2)?

OpenStudy (anonymous):

x – y + 2z = 4 , 3x – 2y + 6z = 12 , 2x – 2y + z = 2 , or 4x – 4y + z = 4

OpenStudy (whpalmer4):

This is easily found by just plugging one of the intercept points into each equation and discarding the equations where you come up with something nonsensical. For example, using point \(X(1,0,0)\): \[x-y+2z=4\]\[1-0+2(0) =4\]Next! The correct answer will produce true equations with all three points.